I need help solving for these two problems! Thank you!

what is the magnitude of the electric field 20 cm away from a 100 nc charge? Show your calculation. What is the electric potential 20 cm away from a 100 nC charge? Show your calculation.
Answer #1

The electric field can be found out using the formula

E = \frac{1}{4\pi\epsilon_0}\frac{q}{r^2} = 9 \times 10^9 \times \frac{100 \times 10^{-9}}{(20 \times 10^{-2})^2} = 22500N/C

The potential is given by the formula

V = \frac{1}{4\pi\epsilon_0}\frac{q}{r} = 9 \times 10^9 \times \frac{100 \times 10^{-9}}{20 \times 10^{-2}} = 4500V

Please upvote if the answer is useful. Comment for doubts and clarifications.